OpenStudy (anonymous):

Please help me! x * 5^3 / 5*6^2 = y what is y if x is 2

OpenStudy (anonymous):

first eliminate the powers x*125/5*36=y Simplify further 125x/180=y now substitute value of x as 2 125(2)/180=y y=250/180 y=1.39 to 2d.p